OpenStudy (anonymous):

A 3.5 µC charge is uniformly distributed on a ring of radius 10.5 cm. Find the electric field on the axis at 1.5 cm from the center of the ring. The value of the Coulomb constant is 8.99 × 10 9 N · m2 /C 2 . Answer in units of N/C

OpenStudy (jamesj):

Step 1: draw a diagram. Now how far is all the charge on the ring from the position 1.5 cm from the center of the ring? That's the distance r you need to use in the Coulomb law calculation

OpenStudy (anonymous):

so 10.5-1.5=9 thats the radius i need to use

OpenStudy (jamesj):

\[ E = \frac{F}{q_2} = \frac{kq_1q_2}{r^2}\frac{1}{q_2} = \frac{kq_1}{r^2} \] You know \( q_1 \) and \( k \), and now hopefully you know r. No ... what you've written there isn't the distance r. Draw a diagram. You will need Pythagorus Theorem.

OpenStudy (anonymous):

i just don't get it

OpenStudy (jamesj):



OpenStudy (jamesj):

The radius of the ring -- the vertical line in this diagram-- is 0.105 m. The distance from the center of the ring -- the horizontal line in this diagram -- is 0.015 m Hence the distance of the point of the triangle on the right is by Pythagorus Theorem...

OpenStudy (jamesj):

\[ \sqrt{(0.105)^2 + (0.015)^2} \ meters \]

OpenStudy (anonymous):

soo.. 3.5 *10^-6(8.99*10^9)/(.106067)^2

OpenStudy (jamesj):

Looks about right.

OpenStudy (anonymous):

thanks!